### v1

@@ 3.1 Masking Strategy
Rather than imposing a masking policy, [35] proposed a framework in which the masking strategy is learned. A teacher network, seeing the intact image, generates a mask that is used to pre-train the student. The objective is to reconstruct the feature representation of the teacher for the masked tokens, as well as the class token [CLS] that is used in generating the mask. The teacher’s parameters are updated using an exponential moving average of the student’s weights. The experiments demonstrated the superiority of the method over other masking strategies.
To boost the performance of any knowledge distillation framework based on feature learning, [36] proposed an auxiliary task. A proportion of the latent feature maps of the student is masked, and a projection layer tries to recover the masked feature maps and match them with those of the teacher. This results in higher performance for a wide range of computer vision tasks.
SemMAE [37] deployed an additional stage before masking. This stage, called semantic part learning, is responsible for learning attention maps that correspond to meaningful semantic components in the image. The training of this stage is performed by embedding the class token provided by a ViT encoder into part embeddings. The resulting embeddings together with the patch embeddings provided by the same encoder are used in an attention layer. Finally, the obtained attention maps are processed by a decoder to reconstruct the original image. In the second stage, the attention maps are used for part segmentation. The masking varies from masking patches in each part to masking entire parts randomly.
[38] introduced Masked Conditional Video Diffusion (MCVD). MCVD leverages different frame masking strategies to train a video diffusion model for unconditional video synthesis and video interpolation. At training time, the method chooses randomly and independently to mask the future or past frames of a given video. This simple strategy allows the diffusion model to perform four different tasks during inference: future and past frame prediction, unconditional generation, and frame interpolation.
[39] added masked image modeling in the training pipeline of GANs. The masking is based on two methods. The first one is called shifted spatial masking and constitutes random masking of the image. The second one is called balanced spectral masking and randomly masks some spectral bands of the image decomposed in the spectral space. The authors observed that these two strategies are orthogonal and help the adversarial training to become more stable.
The work of [40] pioneered the application of Masked Autoencoders (VideoMAE) to video data, adapting the principles of MAE used in image processing to the temporal domain. In this approach, sequences of frames are masked with a masking ratio of approximately 90%. This notably high ratio is strategically chosen to effectively minimize the issue of information leakage that occurs between closely spaced frames. [41] extended this work and introduced VideoMAEv2, which represents a method of scaling up the original VideoMAE.
[42] presented a masked autoencoder pre-training scheme for vision and language medical data. This model consists of two encoders: one based on ViT for images and one based on BERT for texts. Then, the embeddings of the two modalities are jointly processed with a cross-attention module in order to fuse the information. Finally, each input is reconstructed with a separate decoder: a transformer model for images and a simple multi-layer feed-forward network for texts. While the language decoder receives the output of the last layer of the multimodal module, the vision decoder uses an intermediate latent representation. Another important aspect is that different masking ratios are used for different input types.
[45] introduced a method to fine-tune models, with the goal of achieving a more balanced trade-off between in-distribution (ID) and out-of-distribution (OOD) performance. The authors argue that fine-tuning models on a specific dataset tends to enhance their performance on ID data, but this diminishes their performance on OOD data. Therefore, they proposed an approach that involves masking certain patches within an image and replacing them with content from another image. This modified image is then used to train the model under the supervision of the pre-trained model to recognize the masked image features.
[46] observed that existing image denoising methods often overfit on the type of noise seen during training. Their approach focused on improving the generalization capabilities of this type of models by leveraging masked image modeling. To this end, the proposed method masks randomly chosen pixels from the input image and tokens in the self-attention layers of the transformer architecture. The use of token masking in self-attention layers effectively mimics the unreliability observed in tokens during inference, when the data is compromised by various types of noise. This simulation helps to prepare the model for real-world scenarios, where data quality may be inconsistent.
[47] developed a masking strategy based on the feedback provided by the model being trained. The strategy computes the reconstruction loss for each patch and then selectively masks the patches that are more challenging to reconstruct, as indicated by their higher loss values. This approach ensures that the model focuses on the most difficult aspects of the data during training.
[27] showed that masked image modeling is more effective on 3D scenes when certain points are excluded from masking, the so-called Informative Points. Keeping these points unchanged will help to preserve the geometric structure of the scene after the masking process.
[48] proposed DropMAE, an adaptation of MAE for videos. The main observation of [48] is that it is not enough to mask and reconstruct video patches in order to learn spatio-temporal features. The proposed solution is to drop spatial-attention tokens to guide the model towards looking at the temporal information.
OmniMAE, introduced by [49], is similar to the classic MAE framework, but in this case, the model is trained with both video and image data.
The MixMAE [50] pre-training approach integrated elements of both MAE [1] and SimMIM [15]. This method selects two distinct images from a training dataset and extracts random tokens from each. The tokens are then amalgamated to form a new, hybrid image. This composite image undergoes processing by an encoder. Further, a decoder is employed, aiming to reconstruct the original two images from which the mixed image was derived. During the decoding phase, the input provided to the decoder is unmixed and the patches corresponding to the missing tokens in the hybrid image are filled with masked tokens.
[51] proposed SMAUG, which stands out as an efficient framework for video-language pre-training, surpassing previous methodologies. In its pre-training process, SMAUG employed a strategy of masking a significant portion of both frame patches and text tokens, enabling simultaneous masked video modeling and masked language modeling.
[52] argued that, for successful application of MAE in videos, it is crucial to consistently mask video segments across time. Without this consistency, there is a risk of temporal information leakage, rendering the learning task overly simplistic. To address this challenge, they introduced a novel solution that leverages optical flow techniques to generate time-coherent masking volumes. These volumes are then utilized to selectively sample visible tokens, ensuring that the masking process maintains temporal integrity and effectively prevents information leakage.
[30] introduced masked relation modeling to improve self-supervised pre-training on medical images. This masking strategy uses the self-attention mechanism to identify strong dependent regions in the input image and breaks such relations by masking the most important patches for a given patch. On the same note, the cross-attention is applied between images and genome features, aiming to capture the correspondence between these two modalities.
[53] extended the MAE approach to temporal skeleton sequences. Rather than opting for the straightforward method of reconstructing the skeleton sequence directly, this research reconstructed the temporal motion embedded within the sequence, using the masked skeleton sequences as input. Additionally, it used the same motion for guiding the masking of the skeletons.
[54] presented a self-supervised masking-based method for the multi-agent reinforcement learning (RL) setting. For a given timestamp, the observations of all the agents are taken. A part of them are masked and replaced by the values of the previous timestamp, eventually being encoded in a latent space. Then, a reconstruction model is used to generate the original agents’ feature representations from the masked sequence. The authors assert that the resulting feature space is stimulated to learn more about the interaction between agents, while being more temporal aware.
Rather than using a random masking strategy for medical images during MIM pre-training, [55] introduced MedIM, a method that guides the masking based on radiology reports. Firstly, both inputs (the medical image and the corresponding report) are encoded using separate encoders. Then, the text embeddings are split into two subsets, according to the original word categories: MeSH and Sentence tokens. Each subset of text embeddings, together with the visual embeddings, are used to generate a mask that will hide different information. The loss adds the reconstruction errors between the two resulting decoded masked embeddings (with separate heads) and the original image.
[56] presented a masking strategy that is applied to superpixels (contiguous groups of pixels with similar properties), demonstrating its capability on medical image segmentation of skin lesions. Nevertheless, after masking a proportion of superpixels, the same MIM methodology is followed: reconstructing the masked superpixels based on the visible ones. Initially, a base policy [57] is adopted to generate and mask superpixels, after which, the policy is optimized in a self-supervised manner, the model being further pre-trained with the new policy.
Given the scarcity of data containing dental panoramic radiographs, [58] adopted a masked image modeling strategy (either SimMIM or UM-MAE) for pre-training a SwinViT in order to ameliorate its need for a large training dataset. Then, the encoder of the pre-trained SwinViT is taken and used as a backbone in the detection and image segmentation downstream tasks.
[60] proposed a guided masking strategy for MAE, rather than a random policy. Periodically, all images are passed through the encoder and their latent feature representation of the [CLS] token from the last attention layer is extracted in order to compute an importance map of all patches. Using this, the masked patches are sampled (higher chance for more salient patches) and then estimated via reconstruction, while a portion of the least important patches are completely put aside (i.e. not fed to the encoder).
[61] extended the MAE framework to learn temporal feature representations between the frames of a videoclip. Two frames are randomly sampled, the earlier one being split into patches, while the future one is both split and masked. Both images are then separately encoded using Siamese encoders. The resulting embeddings are passed through a decoder with cross-attention (the queries consisting of the future frame’s visible token embeddings and mask tokens) in order to reconstruct the future frame.
Rather than masking at the pixel or patch level, [62] proposed to apply the masks in the frequency domain. The images are converted with fast Fourier transform, then either low or high frequencies are masked, and mapped back into the pixel space. Using an encoder-decoder model, the original image is estimated and transformed into the frequency domain. The objective is to identify the frequencies that were masked.
3.2 Target Features
The early popular frameworks either consist of reconstructing patches of the original image [1, 15] or comparing high-level latent features [17]. Subsequent works adopted different reconstruction targets that are more suitable for the downstream tasks [63, 64], or demonstrated to learn richer feature representations [65]. Finally, MIM has been employed for input signals other than image pixels, and some of these studies involved estimating a different feature type rather than the raw input signal, such as the statistics of a point cloud [66].
Figure 7: SdAE pipeline as presented by [17].
[17] argued that reconstructing the image pixel space does not force the network to learn an ideal representation of the data. Thus, they introduced a self-distillation framework, shown in Figure 7, in which the student branch follows the original MAE flow, while the teacher network (which is not updated by gradient descent, but rather from the weights of the student) only takes the masked patches. The objective is to match the high-level features between the two networks.
[67] proposed the masked image modeling pre-training for 3D meshes. Inspired by ViT, they begin by grouping together faces (each face being represented by a 10D vector) into a patch. Then, they adopt the transformer architecture, using the 3D location of the center of each patch for computing the positional embedding. The algorithm follows MAE: patches are randomly masked with a high masking ratio, the visible patches are passed through the encoder, then the resulting latent embeddings are concatenated with masked tokens (associated with the masked patches), and subsequently decoded. The objective is not only to reconstruct the masked patches (by predicting the coordinates of the vertices), but also the faces of the patch (the 10D representations).
Inspired by MAE, [68] presented a self-supervised masking pre-training strategy that involves Siamese networks. Given a source image and applying transformations on it, two different views (anchor and target) are generated. Then, only the anchor’s tokenized patches are masked, and using the Siamese networks (which follow the ViT architecture), both views are encoded. The latent representations are compared with a set of learnable prototypes in order to generate a distribution, the final goal being that of obtaining matching distributions (predictions of anchor and target). While the anchor’s model is updated using gradient descent, the target’s network parameters are computed as an exponential moving average from the anchor network’s weights. The authors attest that the method greatly improves the performance in few-shot scenarios.
Given a labeled (source) and an unlabeled (target) 3D Point Cloud dataset, the aim of [69] is to transfer the knowledge from the latter to the former by embedding information about common features in an encoder. This model is trained simultaneously on both datasets, but with different objectives. On the one hand, the source dataset is trained in a supervised setting on a specific task. On the other hand, points from the target dataset are randomly masked in arbitrary areas and the objective is to estimate their cardinality (number of points in the neighborhood), position and normal vectors. Therefore, the model benefits from unlabeled data by encapsulating information about the structure of objects.
[70] proposed a video representation learning method based on masking that reconstructs trajectory features rather than static information (like frames). These target features are carefully designed to capture long-term motion changes (object shape and position).
[71] combined image-text contrastive learning (CLIP) [72] and MIM. Their main contribution over the naive approach for combining the two methods consists in using the language space as target space for the reconstruction objective. This is motivated by the intuition that the language features serve as rich semantic representations of the visual signal.
[65] introduced an efficient method for knowledge distillation using the MAE framework. Their approach reconstructs masked patches, while training the student model to replicate the early (low-level) feature maps of the teacher. The efficiency stems from utilizing the MAE framework and from the partial evaluation of the teacher network, as it only requires early feature maps for training the student. This significantly reduces computational cost and training time.
Mask3D [64] enhanced the 2D feature representations of the ViT backbone [16] by integrating 3D priors into the training pipeline. Mask3D utilized RGB-D data within a self-supervised framework, where both color images and corresponding dense depth maps are masked and processed through dual encoders. These encoders project the data into a higher-dimensional space, facilitating a decoder to accurately reconstruct the dense depth map. This method enriches the capability of ViT to handle spatial depth alongside traditional 2D data.
EVA [63] is a ViT model trained to reconstruct CLIP features conditioned on masked image tokens. [63] showed that this self-supervised task is suitable for large scale representation learning. The EVA model scales to one billion parameters using tens of millions of samples, showcasing its potential for handling extensive datasets and complex learning tasks.
Inspired by MIM, [73] introduced a novel approach for boosting the performance of ViT. A number of patches from the original image are shuffled and have their positional encodings masked. Besides the original loss of the downstream task, another objective that tries to predict the masked positional encodings is employed.
GeoMAE [66] is an adaptation of the MAE framework to point clouds. This method changed the usual reconstruction objective and replaced it with centroid prediction, normal estimation and curvature prediction. This change showed significant improvements in downstream tasks such as object detection, segmentation and multi-object tracking.
[74] conducted an empirical analysis on optimal target features for video-language pre-training. They found that spatially-focused image features, extracted using a Swin-B [31] transformer, yield the best results. Additionally, their research incorporates the usual tasks employed in video-language pre-training, such as masked video modeling, masked language modeling and video-text matching.
For their video compression method, [75] employed a transformer-based entropy that is pre-trained using masked image modeling. Some tokens from the current frame are randomly masked, and the model tries to estimate their probability mass functions. More prior information about the last decoded frames is supplied as keys and values. At inference, the prediction is done as an iterative process.
3.3 Objective Function
The primary objective function of MIM is to reconstruct the masked pieces of information from the input signal. Thus, any paper that contributed in this regard, either by improving the reconstruction process or taking a different approach, is included in the category of reconstruction-based frameworks. Inspired by the reconstruction-based techniques, another approach has recently emerged, which applies a contrastive objective function. Furthermore, MIM was integrated in methods that showed great potential in leveraging out-of-distribution unlabeled data [76, 77], and even in adapting a model to the test data distribution [18, 78].
Inspired by the MAE, [79] proposed a similar pre-training framework for 3D Point Clouds, but substituted the reconstruction task with discrimination. A small proportion of points are left unmasked, and subsequently encoded. Then, a subset of the masked ones is sampled (real), along with some random 3D points from the space (fake), and the decoder’s objective is to discriminate between the two. The encoded unmasked points are used in the cross-attention blocks of the decoder. Experiments are conducted for various downstream tasks, in which the method shows considerable performance improvements.
[80] studied the effectiveness of combining MAE and CLIP in a single framework. The conclusion is that the combination brings some benefit when training on a small dataset (tens of millions of samples), but this benefit is smaller or near zero when the experiments are carried out on a much larger dataset (1 billion samples).
Figure 8: CMAE pipeline as presented by [21].
To boost the performance of models that have zero-shot classifying capabilities (such as CLIP), [82] proposed a framework composed of three tasks, from which two are based on MIM. Besides the reconstruction loss of the masked patches, another objective is to minimize the distance between the resulting embeddings of the masked tokens and the embedding of the prepended [CLS] token in a shared projected space.
[20] formally introduced the self-distillation masked autoencoder framework. An input image is divided into patches, some of which are randomly masked. Two encoder-decoder networks (teacher and student) are used to reconstruct the original image. A new objective is employed to minimize the distance between the predictions of the teacher and the student. While the student is trained using gradient descent, the teacher’s weights are computed as an exponential moving average of the student’s.
[83] integrated a masked autoencoder model in a reinforcement learning setting in order to obtain a reward model for exploration. The autoencoder is trained by masking some states from a trajectory and then estimating them. Given a sampled trajectory, for each timestamp, a fixed number of previous states are kept and encoded. Some of the resulting embeddings are then masked, while the rest are passed through the decoder to predict the missing states. In the end, the exploration reward is given by the prediction error.
In their work, [21] combine the reconstruction and contrastive methodologies into one. As presented in Figure 8, they achieve this by using two branches, one for each strategy: the first one with an encoder and pixel decoder being updated at every step, while the second one employs another encoder, which is updated as an exponential moving average from the other encoder, as well as a projection layer and a feature decoder that have the same output vector space. Two views that have a different a shift are created from the input image, one being passed through the first branch (as well as being masked), while the other is fed into the second branch (unaltered). The first branch follows the original MAE framework, with the reconstruction objective. The second branch applies a contrastive loss between the projected embedded features (with the latter encoder) of the second view and the decoded embedded features (i.e. from the former encoder) of the first view.
Figure 9: First step of AMAE pipeline as presented by [84].
Figure 10: Second step of AMAE pipeline as presented by [84].
3.4 Downstream Task
Given the promising results of MIM and its ability to diminish the negative effects on low quantities of data, the pre-training strategy was eagerly applied in many visual tasks and related domains, ranging from image classification [85, 26] and image generation [86, 87] to 3D point clouds [88], graphs [89], and even medical data [84, 90, 91].
MaskGIT [87] trained a generative transformer to reconstruct randomly masked image patches. At inference time, it starts by predicting all the patches and keeps the most confident ones for the next iteration when the rest of the patches are again masked and regenerated. The process continues for a few iterations. Overall the pipeline has actually two stages, the first one encodes the patches into visual tokens with a VQ-Encoder, and the second stage (decoder) receives masked tokens for reconstruction.
[92] proposed to use MAE on videos. The video is split into equally-sized patches that do not overlap along any dimension (i.e. including time), as well as consisting only of two timesteps. As hypothesized and demonstrated by the authors, a high masking ratio (about 90% of the whole input video, unaware of any axis) is used due to redundant data when decoding. Positional (i.e. height and width) and temporal (i.e. time) embeddings are added to the input tokens. The architectures of both the encoder and the decoder are based on ViT. The method follows the same logic as MAE: encoding only the visible tokens, then decoding the complete set tokens.
By formulating the search of a neural network architecture as a graph, a model can be trained and evaluated to estimate its performance. [89] applied a pre-training strategy by masking vertices and then reconstructing the graph (employing an encoder-decoder model). Then, the encoder is fine-tuned to predict the performance of the architecture. This results in higher generalization, while also requiring less data.
MAGVIT [93] trained a transformer-based model for conditional video generation, handling tasks like frame prediction, interpolation, and central outpainting. The model processes videos by selecting a task, and creating conditional tokens from the preprocessed video. The conditional tokens, along with masked and original tokens, form the input used to train the model, which is optimized with three objectives: refining conditional tokens, predicting masked tokens, and reconstructing original tokens. Inference is conducted autoregressively, reducing the masking ratio progressively.
Masked Autoencoder Guided Segmentation at Pixel Resolution (MAESTER) [94] is a masked image modeling approach for the segmentation of cellular images. MAESTER incorporates the masked autoencoder in the training pipeline of a visual transformer to learn token representations relevant to the sub-cellular structure segmentation task (i.e. texture) and performs the segmentation at inference time by clustering the learned representations.
[95] used masked image modeling to learn a latent space from fMRI inputs. After this stage, the authors used the learned latent representations to condition a diffusion model that is able to generate visualizations of the initial visual stimuli.
[96] explored the effectiveness of a pre-trained Vision Transformer (ViT) on masked image modeling tasks within the context of object detection.
Geometry Enhanced Masked Image Modeling (GeoMIM) [97] tackled the problem of 3D detection using multi-view cameras. This method involves pre-training a student network by utilizing masked inputs from multiple camera views. The objective is for the student network to reconstruct the bird’s-eye-view features, leveraging the guidance of a pre-trained LiDAR-based model. This strategy bridges the gap between multi-view camera inputs and LiDAR precision, enhancing the student network’s ability to accurately interpret and reconstruct 3D environments.
[84] leveraged a MAE for detecting anomalies in chest X-rays. The first stage of their pipeline (illustrated in Figure 9) consists of initially pre-training a masked autoencoder. A classification head is then attached to the encoder (whose weights are frozen) to classify normal and abnormal samples, the latter being artificially created. During the second stage, unlabeled data is classified using the model from the previous stage, and the examples that have high confidence are kept. The last step is to employ two different copies of the pre-trained autoencoder, one for each class, and train them separately on the masked reconstruction task. At inference, multiple reconstructions are generated from both autoencoders, an anomalous prediction being detected by a large difference between the mean reconstructed images of the two modules. The steps in the second stage are presented in Figure 10.
In an effort to boost the performance of computer vision models in the medical domain, [98] used visual transformers for multi-label classification of chest X-rays. The large quantities of data required for training a ViT was overcome by pre-training the model on the masked modeling task. Furthermore, the authors proved the superior performance of their method compared with CNN-based architectures.
Similar to [98], [99] demonstrated the benefits of masked image modeling on 3D medical images. They adopted two strategies (original MAE and SimMIM) with various masking hyperparameters, showcasing improved results on two different segmentation tasks.
[91] introduced masked modeling as a pre-training strategy for medical radiography tasks by using the associated radiology reports. While the images are downsampled by half and their unmasked patches are encoded, the text reports are tokenized, randomly masked and then embedded through a look-up table. A global average pooling over the resulting visual embeddings is added to the unmasked text embeddings, and all of them are decoded to obtain the intact report. The image tokens are fed as well through another decoder in order to reconstruct the radiography at its original resolution.
@@ 3.4 Downstream Task
[100] adopted the MAE pre-training strategy for microscopy images. Their objective was to learn a rich feature representation of the cellular morphology of the data. The extensive experiments demonstrated excellent results for predicting biological relationships.
@@ 3.5 Theoretical Analysis
Besides the applied contributions, some pieces of work take another approach: they theorize about various aspects of MIM and dive deeper into its fundamentals. The papers from this category address aspects such as overall understanding [101, 102], the connection between MIM strategies [26], or present certain drawbacks and how to overcome them [103, 104].
@@
Figure 11: The architecture of SSMCTB as presented by [107].
[108] explored the differences in representations learned by deep models through MIM versus traditional supervised training. They observed that MIM encourages models to focus on local patterns across all layers, whereas supervised training emphasizes these patterns only in the initial layers. Additionally, MIM results in a greater diversity among the attention heads compared with supervised methods, suggesting a more nuanced feature recognition within the model.
@@
[103] studied the adversarial robustness of the transformers pre-trained with MIM. Their first observation is that MAE, in particular, has a lower robustness compared with other methods. Moreover, they found that the robustness is related to the reconstruction target. For example, a model trained to reproduce the pixels of an image is prone to adversarial attacks because its focus is on medium and high-frequency features. To ameliorate this issue, the authors proposed a test-time solution based on visual prompts optimized on the frequency domain. These prompts are then included in the input images through prototype-based prompt selection.

### v2

@@ 3.4 Downstream Task
Several works have focused on various tasks in the medical domain. Masked Autoencoder Guided Segmentation at Pixel Resolution (MAESTER) [110] is a masked image modeling approach for the segmentation of cellular images. MAESTER incorporates the masked autoencoder in the training pipeline of a visual transformer to learn token representations relevant to the sub-cellular structure segmentation task (i.e. texture) and performs the segmentation at inference time by clustering the learned representations. [98] used masked image modeling to learn a latent space from fMRI inputs. After this stage, the authors used the learned latent representations to condition a diffusion model that is able to generate visualizations of the initial visual stimuli. [95] leveraged an MAE for detecting anomalies in chest X-rays. The first stage of their pipeline (illustrated in Figure 12) consists of initially pre-training a masked autoencoder. A classification head is then attached to the encoder (whose weights are frozen) to classify normal and abnormal samples, the latter being artificially created. During the second stage, unlabeled data is classified using the model from the previous stage, and the examples that have high confidence are kept. The last step is to employ two different copies of the pre-trained autoencoder, one for each class, and train them separately on the masked reconstruction task. At inference, multiple reconstructions are generated from both autoencoders, an anomalous prediction being detected by a large difference between the mean reconstructed images of the two modules. The steps in the second stage are presented in Figure 13. In an effort to boost the performance of computer vision models in the medical domain, [100] used visual transformers for multi-label classification of chest X-rays. The large quantity of data required for training a ViT was overcome by pre-training the model on the masked modeling task. Furthermore, the authors proved the superior performance of their method compared with CNN-based architectures. Similar to [100], [99] demonstrated the benefits of masked image modeling on 3D medical images. They adopted two strategies (original MAE and SimMIM) with various masking hyperparameters, showcasing improved results on two different segmentation tasks. [111] adopted the MAE pre-training strategy for microscopy images. Their objective was to learn a rich feature representation of the cellular morphology of the data. The extensive experiments demonstrated excellent results for predicting biological relationships. Leveraging both images and text, [112] introduced masked modeling as a pre-training strategy for medical radiography tasks by using the associated radiology reports. While the images are downsampled by half and their unmasked patches are encoded, the text reports are tokenized, randomly masked and then embedded through a look-up table. A global average pooling over the resulting visual embeddings is added to the unmasked text embeddings, and all of them are decoded to obtain the intact report. The image tokens are fed as well through another decoder in order to reconstruct the radiography at its original resolution.
[96] explored the effectiveness of a pre-trained Vision Transformer (ViT) on masked image modeling tasks within the context of object detection. Geometry Enhanced Masked Image Modeling (GeoMIM) [104] tackled the problem of 3D detection using multi-view cameras. This method involves pre-training a student network by utilizing masked inputs from multiple camera views. The objective is for the student network to reconstruct the bird’s-eye-view features, leveraging the guidance of a pre-trained LiDAR-based model. This strategy bridges the gap between multi-view camera inputs and LiDAR precision, enhancing the student network’s ability to accurately interpret and reconstruct 3D environments.
@@
[114] integrated a secondary masked image modeling task in their pose-guided human image generation training pipeline. The authors opted for a transformer-based denoising architecture, rather than a U-Net. As a result, the masked tokens of the target image are reconstructed not only through the self-attention layers, but also using cross-attention modules with the aggregated tokens of the target pose and the source image.
@@ 3.5 Theoretical Analysis
Besides the applied contributions of MIM, some pieces of work take another approach: they theorize about various aspects of MIM and dive deeper into its fundamentals. The papers from this category address aspects such as overall understanding [115, 116, 117, 118, 119, 120], the connection between different MIM strategies [34], or present certain drawbacks and how to overcome them [121, 122].
@@
Figure 15: SSMCTB architecture, courtesy of [31] (image available under the ArXiv perpetual license agreement). SSMCTB falls into the category of reconstruction-based methods depicted in Figure 2, as it employs the Huber loss between the reconstructed output and the input.
[117] explored the differences in representations learned by deep models through MIM versus traditional supervised training. They observed that MIM encourages models to focus on local patterns across all layers, whereas supervised training emphasizes these patterns only in the initial layers. Additionally, MIM results in a greater diversity among the attention heads compared with supervised methods, suggesting a more nuanced feature recognition within the model. In addition, [121] studied the adversarial robustness of the transformers pre-trained with MIM. Their first observation is that MAE, in particular, has a lower robustness compared with other methods. Moreover, they found that the robustness is related to the reconstruction target. For example, a model trained to reproduce the pixels of an image is prone to adversarial attacks because its focus is on medium and high-frequency features. To ameliorate this issue, the authors proposed a test-time solution based on visual prompts optimized on the frequency domain. These prompts are then included in the input images through prototype-based prompt selection.